OpenStudy (anonymous):

can 24 32 and 40 form a right triangle

Mehek (mehek14):

you have to use the Pythagorean Theorem familiar with it?

Mehek (mehek14):

@ggggggggggttt ?

Mehek (mehek14):

\(\Large{a^2 + b^2 = c^2}\)

Mehek (mehek14):

you will take the 2 smaller numbers and substitute them in a and b and the larger number has to equal the square root of the 2 smaller numbers added together

Mehek (mehek14):

so you have the numbers 24 32 40 find \(\Large{24^2 + 32^2 =}\)

OpenStudy (anonymous):

so is it a right triangle

Mehek (mehek14):

did you check?

OpenStudy (anonymous):

it equals 56

Mehek (mehek14):

24 x 24 = ?

OpenStudy (anonymous):

576

Mehek (mehek14):

32 x 32 = ?

OpenStudy (anonymous):

1024

Mehek (mehek14):

so add them 576 + 1024

OpenStudy (anonymous):

1600

Mehek (mehek14):

find the square root of 1600 \(\Large \sqrt{1600}\)

OpenStudy (anonymous):

400

Mehek (mehek14):

you sure? 400 x 400 = 160000 not 1600

Mehek (mehek14):

check again

OpenStudy (anonymous):

i mean 40

OpenStudy (anonymous):

not 400

Mehek (mehek14):

so is it a right angle triangle?

OpenStudy (anonymous):

yes

Mehek (mehek14):

since \(\Large{24^2 + 32^2 = 40^2}\)

OpenStudy (anonymous):

thank you

Mehek (mehek14):

yw \(\ddot \smile\)

OpenStudy (asnaseer):

NOTE: Since all three numbers are divisible by 8 it would have made your job /slightly/ easier if you first divided them all by 8 to get: 3, 4, 5

Mehek (mehek14):

and \(\Large{3^2 + 4^2 = 5^2}\)